OpenStudy (anonymous):

the temperature at the end of the day was 64 degrees. the temperature has increased twice during the day,once by 8 degrees and another time by 4 degrees which expression shows what the temperature was at the start of the day

OpenStudy (anonymous):

... What part of this problem do you not understand?

OpenStudy (anonymous):

I don't know how to actually start it off where would I start in this problem to get the answer?

OpenStudy (anonymous):

Okay then, why don't we call the initial temperature x? We know that the temperature has increased by 8 and 4 degrees at two different times, right? We also know that the final temperature is 64. Therefore, x + 8 + 4 = 64, right?

OpenStudy (anonymous):

But we want to write an equation that shows what the temperature was at the start of the time, meaning that we should isolate x. Bringing 8+4 to the other side gives us x = 64 - 12 or x = 52

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

yes I got it thank you.